# Length-Aware DETR for Robust Moment Retrieval

### v1

@@ 3.3 Length-Aware Decoder
Figure 5: [Left] Group-DETR [6] employs one-to-many matching, where the same labels are utilized across all groups. [Right] Our length-wise matching is one-to-one, and it operates within each length class. By matching only the predictions and ground truths that belong to the same class, this approach enables the creation of length-wise expert queries.
In our previous analysis, we identified that the model struggles to accurately predict both the center and the length of short moments. To address this issue, we propose a length-aware decoder that conditions the moment length, enabling the model to focus more effectively on center prediction. We categorize moment lengths into distinct classes—such as short, middle, and long—by analyzing a cumulative mAP graph and identifying inflection points as boundaries. (Detailed information can be found in the supplementary materials.) The decoder queries are trained using a length-wise matching approach based on these length categories. This categorization creates length-wise expert queries that better handle the specific characteristics of different moment lengths.
Decoder queries with class-pattern. We define $\mathcal{N}_{c}$ as the number of length classes for assigning roles to decoder queries. Drawing inspiration from Anchor-DETR [36], we interpret pattern in pattern embedding as a length category and create class-pattern embeddings $Q_{c}$:
@@
[…]
$\hat{y}^{(k)}=\{\hat{y}_{i}\>\mid\>i^{th}\ query\in\texttt{class}\ k\},\$ (7)
@@
$y^{(k)}=\{y_{i}\>\mid\>\texttt{length}(y_{i})\>\in\>\texttt{class}\>k\>\}.\$ (8)
@@ 4.1 Experimental Setup
Datasets. We utilized three datasets, QVHighlights [19], Charades-STA [10], and TACoS, [30]) for evaluation. QVHighlights consists of over 10k YouTube videos covering various topics such as everyday activities, travel, social activities, and political activities. It contains moments of various lengths distributed evenly and allows for testing our intended aspects effectively, as multiple moments appear within a single video. Considering the diversity and complexity of the dataset, it covers the most realistic and challenging scenario. Charades-STA focuses on daily indoor activities, comprising 9,848 videos with 16,128 annotated queries. The lengths of moments are mostly below 20 seconds. TACoS primarily features activities in the cooking domain, consisting of 127 videos with 18,818 queries. The video lengths vary from very short to nearly 800 seconds, with most moments being shorter than 30 seconds. Unlike QVHighlights and TACoS, which encompass moments of varying lengths, Charades-STA is predominantly composed of shorter samples (under 30 seconds), thus it does not fully align with the problem we aim to address. Nevertheless, we include it in our evaluation to assess the model’s generalization performance.
@@
Implementation details. We divided the classes based on the point where the change in performance was most significant in the validation results of each baseline model. To achieve this, we plotted the cumulative mAP graph with respect to length and identified the inflection points. The thresholds for class division were then determined by calculating the k-means centers of these inflection points. As a result, we set the thresholds for QVHighlights as [12, 36, 65, $\inf$], for Charades-STA as [5.67, 14, $\inf$], and for TACoS as [10, 19, 38, $\inf$], using UVCOM as baseline.
In ForegroundMix, we set $\varepsilon_{\text{cut}}=5$ for QVHighlights and TACoS and $\varepsilon_{\text{cut}}=10$ for Charades-STA to create shorter moments. In Length-Aware Decoder, the number of queries per class $\mathcal{N}_{q}$ was set to 10.
For a fair comparison, we utilize the same features that previous works used. On QVHighlights and TACoS, the video features are extracted from SlowFast [8] and CLIP visual encoder [29]. On Charades-STA, we use two feature types as in previous works. The first type is the video features from SlowFast and CLIP visual encoder, and text features extracted from the CLIP text encoder. The second type is video features extracted from VGG [31] and text features extracted from GloVe [28].
The model is trained for 200 epochs on all datasets with learning rates 1e-4. The batch size is 32 for QVHighlights, 8 for Charades-STA, and 16 for TACoS, following previous methods. We kept all the baseline parameters.
Table 2: Performance gains of our method on the QVHighlights test set across different moment lengths.
Method Short Middle Long All
@@
+Ours 13.47 15.36 39.57 44.93 46.15 52.75 45.03 46.03
+9.52 +8.38 +2.18 +3.81 +3.29 +5.80 +5.03 +6.19
@@
+Ours 13.40 15.45 41.59 46.42 47.51 53.01 46.76 47.23
+8.44 +7.23 +1.51 +3.15 -0.12 +2.21 +3.06 +4.61
@@
+Ours 12.54 16.36 42.40 48.30 46.08 52.22 46.51 47.93
+7.26 +5.69 +0.59 +3.40 +1.13 +3.85 +2.66 +4.75
Table 3: Performance comparison on QVHighlights test set. † indicates training with additional audio features.
@@
+Ours 64.01 48.90 45.03 64.69 47.23 46.03 40.32 66.99
+2.79 +4.41 +5.03 +2.38 +7.78 +6.19 +1.31 +4.86
@@
+Ours 65.43 50.13 46.76 65.66 47.89 47.23 41.41 65.63
+0.77 +1.17 +3.06 +1.68 +4.16 +4.61 +1.50 +2.21
@@
+Ours 63.94 51.10 46.51 65.65 49.44 47.93 40.39 65.50
+0.39 +3.63 +2.66 +2.28 +6.77 +4.75 +0.65 +1.30
@@
+Ours 61.45 40.22 42.31 28.09 56.16 36.10
+2.20 +3.58 +5.92 +4.77 +1.59 +1.97
Table 5: Performance comparison with baseline(QD-DETR) on QVHighlights val set. FGMix, BGMix, and LAD indicate ForeGround Mix, BackGround Mix and Length-Aware Decoder, respectively.
Method Short Middle Long All
@@
FGMix BGMix LAD Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
4.57 7.77 38.89 43.10 42.62 47.44 61.39 46.18 41.06 61.68 41.57 41.00
7.86 12.21 41.42 45.28 43.45 47.69 64.06 48.97 43.84 64.15 43.96 43.32
6.27 11.07 42.74 46.51 44.55 48.57 63.87 49.03 44.62 63.64 44.23 43.86
6.99 10.26 39.48 44.68 43.34 51.23 60.77 46.71 42.36 62.89 45.47 43.91
7.60 13.35 42.95 47.20 45.23 48.42 63.87 50.19 45.37 63.96 46.21 44.95
12.26 16.44 42.89 47.69 45.78 52.75 64.32 50.90 46.83 65.49 48.92 47.84
@@
+5.93 +7.48 +7.26 +4.40 +6.27 +5.75
@@
+5.61 +7.23 +7.11 +4.39 +8.90 +6.84
@@
+10.65 +11.87 +9.71 +7.87 +10.75 +8.45
4.2 Results
Performance with respect to moment length on QVHighlights. In Table 2, our method significantly improves short-moment performance across all baselines. Specifically, for QD-DETR, the R1 average and mAP average for short moments increased by +9.52% and +8.38%, respectively. Moreover, our approach consistently outperforms all baselines in mAP average across all lengths.
Overall performance on QVHighlights. In Table 3, our method yields significant improvements across all metrics, indicating enhanced overall performance across all baselines. Notably, while our primary objective was to improve the short moment performance in Moment Retrieval (MR) by enhancing feature diversity, we also observed substantial performance gains in Highlight Detection (HD). This demonstrates that enhancing feature diversity is an effective strategy that can positively impact other tasks as well.
Overall performance on Charades-STA and TACoS. As shown in Table 4, our method achieved a substantial performance improvement of +5.92% in R1@0.5 on TACoS, which encompasses moments with a broader range of lengths compared to QVHighlights. This significant enhancement demonstrates the superior generalization capabilities of our length-aware approach.
In contrast to other datasets, Charades-STA contains moments with highly restricted lengths, making it less representative of the problem we aim to address. Nevertheless, we included Charades-STA in our evaluation to ensure the robustness of our approach. The results show that our method improved R1@0.7 by +3.58% when using SlowFast and CLIP features, +1.97% using VGG features.
4.3 Ablation Studies and Discussions
Component analysis. In Table 5, we examined the impact of MomentMix and the Length-Aware Decoder on enhancing performance for short moments, observing overall gains. While each component individually improves performance, their combined application leads to even greater improvements. This suggests that our two components, MomentMix and Length-Aware Decoder, each contribute effectively without redundancy, making their combined use the effective approach for tackling the challenge of short-moment retrieval.
Evaluation in few-shot scenarios. To validate the effectiveness of MomentMix as a data augmentation technique, we conducted experiments using 50%, 20%, and 10% of the training data. As shown in Table 6, our method significantly outperformed the baseline (QD-DETR) with substantial performance gains. Specifically, utilizing only half of the training samples with our augmentation surpassed the baseline performance that employed the entire training dataset. Additionally, even in the extreme scenario of using just 10% of the training samples, our method achieved remarkable improvements of +9.71% in R1 average and +8.45% in mAP. These results indicate that MomentMix effectively generates new training samples by enhancing feature diversity.
Figure 6: A qualitative result on QVHighlights val set. Existing models often fail to accurately distinguish between foreground and background, leading to unsuccessful predictions or missed detections of short moments. In contrast, our model is capable of predicting short moments with greater accuracy and robustness.
@@ 4.3 Ablation Studies and Discussions
5 Limitation and Conclusion
Limitation. While we have incorporated length-awareness in the decoder, further investigation into enhancing the encoder is necessary. Moreover, our novel data augmentation method for MR improves performance but also increases training time. Thus, future research should focus on developing more efficient approaches.
Conclusion. In this study, we addressed the limitations of short-moment retrieval in existing DETR-based approaches from both data and model perspectives. To overcome the data-centric issue of limited feature diversity in short moments, we introduced MomentMix, which leverages two mix-based data augmentation strategies: ForegroundMix and BackgroundMix. These strategies enhance the feature representations of both foreground and background elements. On the model side, we identified inaccuracies in center predictions for short moments and proposed a Length-Aware Decoder with a novel bipartite matching process conditioned on moment length. This approach leverages length expert queries to improve center prediction accuracy. Extensive experiments demonstrate that our method surpasses state-of-the-art DETR-based moment retrieval models in terms of R1 and mAP on benchmark datasets. Furthermore, our methodology can be seamlessly integrated with other DETR-based models, paving the way for future advancements in the field.

### v2

@@ 3.3 Length-Aware Decoder
Figure 5: Overview of the Length-Aware Decoder (LAD). The model applies a length-wise bipartite matching strategy, where predictions and ground truths of the same class are exclusively paired. This facilitates the creation of class-specific decoder queries, enhancing sequence modeling performance.
In our previous analysis, we identified that the model struggles to accurately predict both the center and the length of short moments. To address this issue, we propose Length-Aware Decoder (LAD) that conditions the moment length, enabling the model to focus more effectively on center prediction. We categorize moment lengths into distinct classes—such as short, middle, and long—by analyzing a cumulative mAP graph and identifying inflection points as boundaries. (Detailed information can be found in the supplementary materials.) The decoder queries are trained using a length-wise matching approach based on these length categories. This categorization creates length-wise expert queries that better handle the specific characteristics of different moment lengths.
Decoder Queries with Class-Pattern. We define $\mathcal{N}_{c}$ as the number of length classes for assigning roles to decoder queries. Drawing inspiration from Anchor-DETR [34], we interpret pattern in pattern embedding as a length category and create class-pattern embeddings $Q_{c}$ with dimension $d$:
@@
We denote $\hat{y}=\{\hat{y}_{i}\}^{\mathcal{N}_{c}\mathcal{N}_{q}}_{i=1}$ as all the predicted moments from the decoder head, where $\mathcal{N}_{c}$ and $\mathcal{N}_{q}$ are a number of classes and a number of queries for each class $k\in\texttt{length-classes}$, respectively. Then, the predictions belonging to class $k$ can be denoted as:
$\hat{y}^{(k)}=\{\hat{y}_{i}\>\mid\>i^{th}\ query\in\texttt{class}\ k\},\\$ (7)
@@
$y^{(k)}=\{y_{i}\>\mid\>\texttt{length}(y_{i})\>\in\>\texttt{class}\>k\>\}.\\$ (8)
@@ 4.1 Experimental Setup
Datasets. We utilized three datasets, QVHighlights [17], Charades-STA [8], and TACoS [28]) for evaluation. QVHighlights consists of over 10k YouTube videos covering various topics such as everyday activities, travel, social activities, and political activities. It contains moments of various lengths distributed evenly and allows for testing our intended aspects effectively, as multiple moments appear within a single video. Considering the diversity and complexity of the dataset, it covers the most realistic and challenging scenario. Charades-STA focuses on daily indoor activities, comprising 9,848 videos with 16,128 annotated queries. The lengths of moments are mostly below 20 seconds. TACoS primarily features activities in the cooking domain, consisting of 127 videos with 18,818 queries. The video lengths vary from very short to nearly 800 seconds, with most moments being shorter than 30 seconds.
@@
Implementation Details. We divided the classes based on the point where the change in performance was most significant in the validation results of each baseline model. To achieve this, we plotted the cumulative mAP graph with respect to length and identified the inflection points. The thresholds for class division were then determined by calculating the k-means centers of these inflection points. As a result, we set the thresholds for QVHighlights as [12, 36, 65, $\inf$], for Charades-STA as [5.67, 14, $\inf$], and for TACoS as [10, 19, 38, $\inf$], using UVCOM as baseline. In MomentMix, we set $\varepsilon_{\text{cut}}=5$ for QVHighlights and $\varepsilon_{\text{cut}}=10$ for TACoS and Charades-STA. In LAD, the number of queries per class $\mathcal{N}_{q}$ was set to 10. For a fair comparison, we utilize the same features that previous works used. On QVHighlights and TACoS, the video features are extracted from SlowFast [6] and CLIP visual encoder [27]. On Charades-STA, we use two feature types as in previous works. The first type is the video features from SlowFast and CLIP visual encoder, and text features extracted from the CLIP text encoder. The second type is video features extracted from VGG [30] and text features extracted from GloVe [26]. The model is trained for 200 epochs on all datasets with learning rates 1e-4. The batch size is 32 for QVHighlights, 8 for Charades-STA, and 16 for TACoS, following previous methods. We kept all the baseline parameters.
4.2 Results
Table 1: Performance gains of our method on the QVHighlights test set across different moment lengths. ‡ means test results from checkpoint provided by authors.
Method Short Middle Long Full
@@ 4.2 Results
+Ours 10.33 13.79 41.32 46.94 45.40 52.45 45.02 46.61
$^{\textbf{{(+6.38)}}}$ $^{\textbf{{(+6.81)}}}$ $^{\textbf{{(+3.93)}}}$ $^{\textbf{{(+5.82)}}}$ $^{\textbf{{(+2.54)}}}$ $^{\textbf{{(+5.50)}}}$ $^{\textbf{{(+5.01)}}}$ $^{\textbf{{(+6.77)}}}$
@@
+Ours 10.30 14.57 42.54 47.90 47.61 53.58 46.59 47.77
$^{\textbf{{(+5.35)}}}$ $^{\textbf{{(+6.35)}}}$ $^{\textbf{{(+2.45)}}}$ $^{\textbf{{(+4.63)}}}$ $^{\textbf{{(-0.02)}}}$ $^{\textbf{{(+2.78)}}}$ $^{\textbf{{(+2.89)}}}$ $^{\textbf{{(+5.15)}}}$
@@
+Ours 11.58 16.35 43.22 52.62 46.48 52.62 46.92 48.21
$^{\textbf{{(+6.31)}}}$ $^{\textbf{{(+5.68)}}}$ $^{\textbf{{(+1.41)}}}$ $^{\textbf{{(+7.72)}}}$ $^{\textbf{{(+1.53)}}}$ $^{\textbf{{(+4.25)}}}$ $^{\textbf{{(+3.07)}}}$ $^{\textbf{{(+5.03)}}}$
Performance with respect to Moment Length on QVHighlights. In Table 1, our method significantly improves short-moment performance across all baselines. Specifically, for QD-DETR, the R1 average and mAP average for short moments increased by +6.38%p and +6.81%p, respectively. Moreover, our approach consistently outperforms all baselines in mAP average across all lengths.
Overall Performance on QVHighlights. In Table 2, our method yields significant improvements across all metrics, indicating enhanced overall performance across all baselines. Notably, while our primary objective was to improve the short moment performance in Moment Retrieval (MR) by enhancing feature diversity, we also observed substantial performance gains in Highlight Detection (HD). This demonstrates that enhancing feature diversity is an effective strategy that can positively impact other tasks.
Table 2: Performance comparison on QVHighlights test set. † indicates training with additional audio features. ‡ means test results from checkpoint provided by authors.
@@
+Ours 63.62 48.77 45.02 65.50 47.78 46.61 40.35 64.46
$^{\textbf{{(+2.40)}}}$ $^{\textbf{{(+4.28)}}}$ $^{\textbf{{(+5.01)}}}$ $^{\textbf{{(+3.19)}}}$ $^{\textbf{{(+8.33)}}}$ $^{\textbf{{(+6.77)}}}$ $^{\textbf{{(+1.34)}}}$ $^{\textbf{{(+2.33)}}}$
@@
+Ours 65.63 51.23 46.59 66.89 49.04 47.77 41.54 66.02
$^{\textbf{{(+0.97)}}}$ $^{\textbf{{(+2.27)}}}$ $^{\textbf{{(+2.89)}}}$ $^{\textbf{{(+2.91)}}}$ $^{\textbf{{(+5.31)}}}$ $^{\textbf{{(+5.15)}}}$ $^{\textbf{{(+1.63)}}}$ $^{\textbf{{(+2.60)}}}$
@@
+Ours 65.37 50.71 46.92 66.65 49.22 48.21 40.91 66.54
$^{\textbf{{(+1.82)}}}$ $^{\textbf{{(+3.24)}}}$ $^{\textbf{{(+3.07)}}}$ $^{\textbf{{(+3.28)}}}$ $^{\textbf{{(+6.55)}}}$ $^{\textbf{{(+5.03)}}}$ $^{\textbf{{(+1.17)}}}$ $^{\textbf{{(+2.34)}}}$
Overall Performance on Charades-STA and TACoS. As shown in Table 3, our method significantly improves performance on Charades-STA, increasing R1@0.7 by +3.58%p with SlowFast and CLIP features, and by +1.97%p with VGG features. Furthermore, our method achieves notable gains (+5.82%p in R1@0.5) on TACoS, which encompasses moments with a broader range of lengths compared to QVHighlights. This substantial improvement demonstrates the strong generalization capability of our length-aware approach.
@@
+Ours 61.45 40.22 42.21 28.02 56.16 36.10
$^{\textbf{{(+2.20)}}}$ $^{\textbf{{(+3.58)}}}$ $^{\textbf{{(+5.82)}}}$ $^{\textbf{{(+4.70)}}}$ $^{\textbf{{(+1.59)}}}$ $^{\textbf{{(+1.97)}}}$
4.3 Ablation Studies and Discussions
Component Analysis. In Table 4, we examined the impact of MomentMix and the Length-Aware Decoder on enhancing performance for short moments, observing overall gains. While each component individually improves performance, their combined application leads to even greater improvements. This suggests that our two components, MomentMix and LAD, each contribute effectively without redundancy, making their combined use the effective approach for tackling the challenge of short-moment retrieval.
Table 4: Performance comparison with baseline(QD-DETR) on QVHighlights val set. MMix, and LAD indicate MomentMix and Length-Aware Decoder, respectively.
Short Middle Long Full
MMix LAD R1 mAP R1 mAP R1 mAP R1 mAP
4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
6.48 11.44 42.96 46.88 44.15 49.57 44.66 44.68
8.76 11.01 40.55 45.53 43.69 50.76 43.65 44.48
11.07 15.27 43.12 48.53 44.39 52.65 46.13 47.70
Table 5: Performance comparison of MomentMix and other augmentations—Gaussian Noise, Random Drop, Random Crop, and Random Insert—on the QVHighlights val set with QD-DETR as the baseline. Other augmentations fail to improve performance, while MomentMix achieves significant gains in short moments and overall robustness in moment retrieval tasks.
Method Short Middle Long Full
@@ 4.3 Ablation Studies and Discussions
Baseline 4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
Gaussian Noise 5.20 7.62 37.97 42.64 43.54 48.18 41.01 40.58
Random Drop 4.15 7.24 39.87 44.05 42.44 47.24 41.48 40.95
Random Crop [20] 5.31 8.99 39.34 43.03 41.01 46.00 40.95 41.00
Random Insert [20] 6.71 8.72 38.59 42.77 41.69 45.94 41.12 40.46
MomentMix 6.48 11.44 42.96 46.88 44.15 49.57 44.66 44.68
Effect of MomentMix. To validate the effectiveness of MomentMix, a novel mixing-based augmentation for moment retrieval (MR) for short moment, we compare it with four baseline augmentations: Gaussian Noise, Random Drop, Random Crop and Random Insert. For Gaussian Noise and Random Drop, we report only the best-performing hyperparameters: adding Gaussian noise with a standard deviation of 0.01 and masking 50% of frame features, respectively. Random Crop and Random Insert were originally proposed for proposal-based long-video MR and adapted by us for proposal-free DETR-based methods. In contrast, our proposed MomentMix uniquely employs foreground mixing (ForegroundMix) and background replacement (BackgroundMix). Specifically, ForegroundMix cuts and mixes foreground segments rather than simply cropping them, while BackgroundMix replaces each background segment independently with clips from different videos, instead of inserting a single foreground segment. As shown in Table 5, MomentMix consistently outperforms baseline augmentations, significantly boosting performance on short moments through enhanced feature diversity (see Fig. A1 in Supple.), while also improving longer moments.
@@
$^{\textbf{{(+5.93)}}}$ $^{\textbf{{(+7.48)}}}$ $^{\textbf{{(+7.26)}}}$ $^{\textbf{{(+4.40)}}}$ $^{\textbf{{(+6.27)}}}$ $^{\textbf{{(+5.75)}}}$
@@
$^{\textbf{{(+5.61)}}}$ $^{\textbf{{(+7.23)}}}$ $^{\textbf{{(+7.11)}}}$ $^{\textbf{{(+4.39)}}}$ $^{\textbf{{(+8.90)}}}$ $^{\textbf{{(+6.84)}}}$
@@
$^{\textbf{{(+10.65)}}}$ $^{\textbf{{(+11.87)}}}$ $^{\textbf{{(+9.71)}}}$ $^{\textbf{{(+7.87)}}}$ $^{\textbf{{(+10.75)}}}$ $^{\textbf{{(+8.45)}}}$
Evaluation in Few-shot Scenarios. To validate the effectiveness of MomentMix as a data augmentation technique, we conducted experiments using 50%, 20%, and 10% of the training data. As shown in Table 6, our method outperformed the baseline (QD-DETR) with substantial performance gains. Specifically, utilizing only half of the training samples with our method surpassed the baseline performance that used the entire dataset. Additionally, even in the extreme scenario of using just 10% of the training samples, our method achieved remarkable improvements of +9.71%p in the R1 average and +8.45%p in the mAP average. These results indicate that MomentMix effectively generates new training samples by enhancing feature diversity.
Attention in LAD. To examine whether our LAD attends to the characteristics of each length-specific query as intended, we investigated attention between these queries and encoder features. In Fig. 6, the existing method [25] predicts moments by focusing on the boundary regardless of length, which we refer to as boundary attention of the moment. However, when applying our method, we observed that attention varies according to the characteristics of each length-specific query. For short moments, attention is focused more on the inside of the moment rather than the boundary, which we refer to as inside attention of the moment. On the other side, for very long moments, attention is directed toward the boundary similar to the existing method. This indicates that our query attends to the length rather than the center of the moment appears to be more efficient for detection, hence the boundary attention. Conversely, for short moments, focusing on the center rather than the length is more efficient, resulting in a moment of inside attention.
Figure 6: Decoder query attention to encoder features on QVHighlights. The top figure shows baseline attention, while the bottom shows attention after applying LAD. Length-expert queries are highlighted in red text. LAD effectively aligns attention patterns with the corresponding length characteristics.
Figure 7: A qualitative result on QVHighlights. Existing models often fail to accurately distinguish between foreground and background, leading to unsuccessful predictions or missed detections of short moments. In contrast, our model is capable of predicting short moments with greater accuracy and robustness.
@@
5 Discussion
Conclusion. This study addressed the limitations of short-moment retrieval in existing DETR-based approaches from both data and model perspectives. To mitigate the issue of limited feature diversity in short moments, we introduced MomentMix, a novel two-stage mix-based data augmentation. These strategies enhance the feature representations of both foreground and background elements for short-moment data samples. On the model side, we identified inaccuracies in center predictions for short moments and proposed a Length-Aware Decoder with a novel bipartite matching process conditioned on moment length. This approach leverages length expert queries to improve center prediction accuracy. Extensive experiments demonstrate that our method outperforms state-of-the-art DETR-based moment retrieval models in terms of R1 and mAP on benchmark datasets. Moreover, our methodology can be seamlessly integrated with other DETR-based models, paving the way for future advancements in the field.
Future Work. In future work, we plan to explore adapting and extending our augmentation approach to non-DETR architectures. Additionally, conducting further experiments on larger-scale models could provide deeper insights into the scalability and robustness of our method.